OpenStudy (anonymous):

SIMPLIFY : square root of 27x^8y^5

OpenStudy (vijay):

\[3y^2 x^4 \sqrt{3y}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

can you explain how you got that please!

OpenStudy (vijay):

\[\sqrt{3*3^2*(x^4)^2*(y^2)^2*y}\]

OpenStudy (anonymous):

in words.... please

OpenStudy (anonymous):

from your final answer in words can you please explain what you did.. i am looking for explinations.

OpenStudy (vijay):

After that , when you square root them , power 2 will ignore by sqrt. then 3 and y will be inside the sqrt ., others will come out.

OpenStudy (anonymous):

no i mean from the begining please!

OpenStudy (anonymous):

so looking at the equation what is the first step that i would taking in simplifying it?

OpenStudy (preetha):

I think what Vijay did was to break each term down into something x a square. So 27 is 3 x (3x3) or 3 x 3^2. x^8 is x^4 squared. And y^5 is y^4 x y.

OpenStudy (vijay):

Thanks preetha

OpenStudy (preetha):

Then you take the square root of this equation he has written out so nicely with the equation editor. So if you have 3 x (3x3) the sq. root of that is sqrt 3 (3), sqrt of (x4)^2 is x^4 and sqrt of y(y^4) is sqrt y (y^2).

OpenStudy (preetha):

Then you put the three terms together. Hope that helps. I am just too lazy to use the equation editor. Good job Vijay.

OpenStudy (anonymous):

hello?

OpenStudy (vijay):

yes

OpenStudy (anonymous):

nevermind i got it ! thanks. but here is one : twelve network printers can produce a set of documents in 20 hours. how long would it take 15 printers to get the same job done? (the time is inversly proportional tot the # of printers used for job)

OpenStudy (vijay):

=(20/12) * 15 =25 hours

OpenStudy (anonymous):

why did you multiply by 15?

OpenStudy (anonymous):

the answer says 16 hrs... i am unsure how the book got that.

OpenStudy (vijay):

yes, 25 is wrong, wait i l do it

OpenStudy (anonymous):

i got it. 15/12 = 20/x 15x= 12*20 15/240 = 16

OpenStudy (vijay):

yes, gud

OpenStudy (anonymous):

hey i just posted something do you know how to factor?